Homework - Chapter 2 Saved 5 Exercise 2-10 Applying Overhead Cost to a Job [LO2-2] 14.28 points Sigma Corporation applies overhead cost to jobs on the basis of direct labor cost. Job V, which was started and completed during the current period, shows charges of $5,200 for direct materials, $8,700 for direct labor, and $5,655 for overhead on its job cost sheet Job W, which is still in process at year-end, shows charges of $2,900 for direct materials and $4,100 for direct labor. Required: 1a. Should any overhead cost be applied to Job W at year-end? eBook Yes O No Print References b. How much overhead cost should be applied to Job W? Overhead cost i. This is a numeric cell, so please enter numbers only. 2. How will the costs included in Job Ws job cost sheet be reported within Sigma Corporations financial statements at the end of the year? O Raw Materials Work-in-Process O Finished Goods
Answer #1

1a) Yes, overhead should be applied on Job W

1b) Overhead rate = 5655/8700 = 65% of direct labor cost

Overhead cost on job W = 4100*65% = $2665

2) Job W should be included in financial statement as work in process

So answer is b) Work in process